A cardiac defibrillator is used to shock a heart that is beating erratically. A capacitor in this device is charged to 5.0 kV k V and stores 1100 J J of energy.

capacitance?
Answer #1

Given,

Potential difference ,V = 5000

Energy of the capacitor, E = 1/2*C*V² =1100 J

0.5*C*(5000)² = 1100 J

C = 2200/(5000)²

C = 0.000088 F = 88 microfarad is the correct answer